OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation x - 2y = 4 using the x- and y- intercepts method ?

OpenStudy (anonymous):

What are your options?

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

Plot the y-intercept (0, -2). Plot the intercept (4, 0). Count up 1 and right 2 from the y-intercept. Draw a line through the two points.

OpenStudy (anonymous):

those are the options , can you help me

OpenStudy (anonymous):

Ok I remember now!:) This is x and y intercept method: Make x=0 to find y-intercept. x - 2y = 4 0 - 2y = 4 -2y = 4 /-2 y = -2 (0, -2) Make y=0 to find x-intercept x - 2y = 4 x - 2(0) = 4 x - 0 = 4 x = 4 (4, 0) And then cross a line through the two points. The 3rd option is the answer.Why? You only do that step when you change the equation to slope-incercept and solve.

OpenStudy (anonymous):

So 3)Count up 1 and right 2 from the y-intercept. Sorry I took so long to realize that!!! x- and y- intercept method: plot x intercept plot y intercept cross a line between points

OpenStudy (anonymous):

thank you i got it now ! :)